OpenStudy (anonymous):

The quotient of a number and -2/3 is -9/10 what is the number?

OpenStudy (jdoe0001):

so $$\huge { \cfrac{n}{-\frac{2}{3}} = -\cfrac{9}{10} \implies \cfrac{\frac{n}{1}}{-\frac{2}{3}}= -\cfrac{9}{10} } $$ simplify the left side, what fraction would that give you?